# Mechanistic Unlearning: Robust Knowledge Unlearning and Editing via Mechanistic Localization

### v1

@@ A.2.3 Mechanism Weight Analysis
This supports our argument that OT methods target high logit-diff extraction mechanisms, rather than the fact lookup mechanisms that enrich the latent stream with the correct attributes, which decreases the robustness of edits/unlearning. It is important to note that since our manual interpretability localization is based on our discovered mechanisms, this does not serve as an evaluation of manual interpretability, but rather only of causal tracing and attribution patching.
A.3 Additional CounterFact Weight Masking Results
We report a comparison of all localizations across discretization thresholds for normal and MCQ forget sets in Figure 17 and Figure 17. We see that manual interpretability outperforms all other methods of localization in preserving maintain accuracy while decreasing forget accuracy.
We perform additional adversarial analysis of accuracies across different discretization thresholds. We report the ”paraphrase” and ”neighborhood” adversarial results in Figure 19 and Figure 19.
A.4 Hyperparameters
For localized fine-tuning, we use $\lambda_{1}=0.2$ for forgetting basketball associations and $\lambda_{1}=1$ for forgetting particular athlete associations. For weight masking on Sports Facts, we use $\lambda_{1}=0.3$ for both basketball and athlete associations. $\lambda_{2}$ and $\lambda_{3}$, the retain and SFT loss coefficients, were both set to 1 across all experiments. For weight masking regularization, we used $\lambda_{4}=1*10^{-7}$ (where our regularization loss was the total sum of absolute weight mask values).
For weight masking on CounterFact, we use $\lambda_{1}=\lambda_{2}$ = $\lambda_{3}$ = 1. For weight masking regularization, we again used $\lambda_{4}=1*10^{-7}$ (where our regularization loss was the total sum of absolute weight mask values).
For localized fine-tuning on Gemma, we trained using 50 iterations of batch size 4 with 16 accumulation steps, using an AdamW optimizer  (Kingma & Ba, 2017) with 0 weight decay, a learning rate of $1*10^{-5}$, and a cosine annealing scheduler for both basketball and athletes.
For weight masking, we unlearned using 50 iterations of batch size 10 with 15 accumulation steps, using an SGD optimizer (for memory efficiency) with learning rate of $1*10^{-2}$ and clamping the mask values between 0 and 1 every update step.
A.5 Evaluation Details
A.5.1 Train-Test Splits
We split the basketball $D_{forget}$ set and both $D_{retain}$ sets (basketball and athletes) into an 80%-20% train-test-split, and all of our reported numbers are on the test set. We do not split the $D_{forget}$ set of 16 athletes, because we wish to test if the model has unlearned the athletes it was trained to unlearn.
A.5.2 Details on Adversarial Relearning
We retrain the model with only two athletes in $D_{forget}$ for multiple iterations (along with a standard retain and SFT loss), in both the sport and athlete unlearning scenarios. In practice, for basketball this looks like giving the model the same batch of only Boris Diaw and Jae Crowder multiple times, and for athletes we give a batch of only DeForest Buckner and Walter Payton. We retrain with a rank-64 LoRA on all linear modules.
A.5.3 Details on Latent Knowledge
We don’t follow the same methodology as Patil et al. (2023) because we only care about the same three possible tokens, so it isn’t applicable to apply their search-budget methodology and we instead try linear probes. In general, we don’t consider linear probing to be a realistic threat model for beating unlearning, as attackers need white-box access and labels for large subsets of the forget set, but we do these tests for an approximate upper bound of accessible information by a capable-enough adversary.
We train three linear probes (Alain & Bengio, 2018) for every model and layer, one for each sport (to predict True or False with a base rate of  66%), on samples from both the forget and retain datasets. We train each probe on both forget and retain samples because for sports, there is only one forget sport (so the answer would be constant if we trained different probes), and for athletes we only have 16 total examples that must further be split into train-test.
For athletes, we split the forget set into a 50%-50% train-test split, so the probe training dataset includes 8 of the forgotten athletes (along with the standard retain train split) and the test set includes the other 8 (along with the retain test split). For sports, we use the standard basketball train and test split. Then, as a measure of aggregated accuracy, we only consider a test sample to be correct if probes for all three sports are correct.

### v2

@@ A.5 Mechanism Weight Analysis
This supports our argument that OT methods target high logit-diff extraction mechanisms, rather than the fact lookup mechanisms that enrich the latent stream with the correct attributes, which decreases the robustness of edits/unlearning. It is important to note that since our FLU localization is based on our discovered mechanisms, this does not serve as an evaluation of FLU (since by definition FLU localization will only localize the FLU mechanism), but rather only of causal tracing and attribution patching.
A.6 Hyperparameters
Across all tasks except Sequential-CounterFact-Editing and all models, we fine tune using 50 iterations of batch size 4 with 16 accumulation steps, using an AdamW optimizer  (Kingma & Ba, 2017) with 0 weight decay and a cosine annealing scheduler. For Gemma-2-9b, we are forced to use an 8-bit optimizer to fit our training in the memory of 1 GPU. We find that the optimal learning rate is quite sensitive to the localization used and the edit task, so we first sweep over learning rates to find reasonable learning rates. We sweep over the learning rates of 2e-6, 5e-6, 1e-5, 2e-5, 5e-5, and 1e-4, training models over 50 iterations with all $\lambda$s set to 1. We also tune the $\lambda_{1}$ parameter associated with the $L_{\text{injection}}$ cross entropy loss. We don’t tune the other $\lambda$ parameters because they are all maintenance losses, and setting them to 1 works sufficiently to maintain performance across almost all setups.
For the Sequential-CounterFact-Editing task, we use the same hyperparameters from the CounterFact-Editing task and we train for 100 total iterations rather than 50, using 25 for each subset of 16 facts. We choose 25 iterations because it balances between being half the number of iterations we typically use per subset of that size, and also double the number of steps overall as we use in CounterFact-Editing.
[…]
A.6.1 Sample Hyperparameter Sweep
In Table 4, we show the full results of one sweep, optimizing learning rate for editing 16 facts from CounterFact. Especially for No Localization, some learning rates fail to edit in the correct answer with high accuracy, or fail to maintain accuracy on the maintain set. In Table 5, we see that editing results are not particularly sensitive to the coefficient used with the injection cross entropy loss.
Table 4: Gemma-7b learning rate sweep, editing 16 CounterFact facts.
Pile Accuracy $\uparrow$ Forget Accuracy $\downarrow$ Edit Accuracy $\uparrow$ Maintained Accuracy $\uparrow$ Overall Score $\uparrow$
FLU
LR 0.0001 0.488 0.000 1.000 0.698 3.186
LR 1e-05 0.513 0.000 1.000 0.975 3.488
LR 2e-05 0.542 0.000 1.000 0.950 3.492
LR 2e-06 0.499 0.007 0.961 0.869 3.323
LR 5e-05 0.520 0.000 1.000 0.822 3.342
LR 5e-06 0.528 0.000 0.999 0.980 3.507
Localized CT
LR 0.0001 0.462 0.001 0.999 0.697 3.157
LR 1e-05 0.513 0.000 1.000 0.984 3.496
LR 2e-05 0.507 0.000 1.000 0.961 3.467
LR 2e-06 0.540 0.036 0.921 0.849 3.274
LR 5e-05 0.488 0.000 1.000 0.846 3.334
LR 5e-06 0.537 0.000 1.000 0.982 3.519
Nonlocalized
LR 0.0001 0.062 0.032 0.557 0.094 1.680
LR 1e-05 0.520 0.000 1.000 0.892 3.412
LR 2e-05 0.479 0.001 0.998 0.807 3.284
LR 2e-06 0.529 0.000 1.000 0.982 3.510
LR 5e-05 0.046 0.034 0.710 0.092 1.815
LR 5e-06 0.536 0.000 1.000 0.988 3.524
Table 5: Gemma-7b inject loss coefficient sweep, editing 16 CounterFact facts.
Pile Accuracy $\uparrow$ Forget Accuracy $\downarrow$ Edit Accuracy $\uparrow$ Maintained Accuracy $\uparrow$ Overall Score $\uparrow$
FLU
FC 0.1 0.538 0.000 0.998 0.990 3.525
FC 0.2 0.510 0.000 0.999 0.981 3.491
FC 0.5 0.508 0.000 0.999 0.985 3.493
FC 1 0.510 0.000 0.999 0.973 3.483
FC 2 0.532 0.000 1.000 0.984 3.515
FC 5 0.534 0.000 1.000 0.985 3.518
Localized CT
FC 0.1 0.532 0.004 0.995 0.983 3.506
FC 0.2 0.535 0.001 0.998 0.986 3.518
FC 0.5 0.524 0.000 0.999 0.974 3.497
FC 1 0.519 0.000 1.000 0.988 3.507
FC 2 0.536 0.000 1.000 0.979 3.514
FC 5 0.518 0.000 1.000 0.975 3.493
Nonlocalized
FC 0.1 0.525 0.001 0.998 0.969 3.492
FC 0.2 0.562 0.000 0.999 0.982 3.543
FC 0.5 0.524 0.000 1.000 0.962 3.486
FC 1 0.531 0.000 1.000 0.980 3.511
FC 2 0.514 0.000 1.000 0.978 3.492
FC 5 0.528 0.000 1.000 0.975 3.503
A.6.2 All Hyperparameters Used
Table 6 has all learning rates used and Table 7 has all injection loss coefficients used.
Table 6: Optimal learning rates for different models, task types, and localizations.
Model 64 athletes to random sport Basketball Athletes to Golf 16 CounterFact facts 64 CounterFact facts
Gemma
FLU $1\times 10^{-5}$ $2\times 10^{-5}$ $5\times 10^{-6}$ $1\times 10^{-5}$
Localized CT $1\times 10^{-5}$ $5\times 10^{-6}$ $5\times 10^{-6}$ $2\times 10^{-5}$
Nonlocalized $2\times 10^{-6}$ $5\times 10^{-6}$ $5\times 10^{-6}$ $2\times 10^{-6}$
Gemma 2
FLU $1\times 10^{-5}$ $5\times 10^{-5}$ $5\times 10^{-5}$ $2\times 10^{-5}$
Localized CT $5\times 10^{-5}$ $5\times 10^{-5}$ $5\times 10^{-5}$ $2\times 10^{-5}$
Nonlocalized $2\times 10^{-5}$ $1\times 10^{-5}$ $5\times 10^{-6}$ $5\times 10^{-6}$
Llama 3
FLU $5\times 10^{-5}$ $5\times 10^{-5}$ $1\times 10^{-4}$ $5\times 10^{-5}$
Localized CT $1\times 10^{-4}$ $5\times 10^{-5}$ $2\times 10^{-5}$ $5\times 10^{-5}$
Nonlocalized $2\times 10^{-5}$ $1\times 10^{-5}$ $2\times 10^{-5}$ $1\times 10^{-5}$
Table 7: Optimal inject loss coefficients for different models, task types, and localizations.
Model 64 athletes to random sport Basketball Athletes to Golf 16 CounterFact facts 64 CounterFact facts
Gemma
FLU 5.0 5.0 0.1 2.0
Localized CT 0.1 1.0 0.2 1.0
Nonlocalized 0.2 1.0 0.2 1.0
Gemma 2
FLU 1.0 5.0 2.0 0.5
Localized CT 5.0 5.0 2.0 2.0
Nonlocalized 5.0 5.0 1.0 1.0
Llama 3
FLU 5.0 0.2 1.0 0.2
Localized CT 2.0 1.0 0.1 2.0
Nonlocalized 2.0 0.1 2.0 0.2
A.7 Evaluation Details
A.7.1 Details on Standard Prompt Evaluations
We report standard metrics of Forget Error, Edit Accuracy, and Maintain Accuracy, in the same prompt format that the models were trained on. These metrics are optimized by the loss, so we expect all localizations to do almost perfectly on these evaluations. Figures 26, 27, 28 and 29 show that localizations perform approximately equivalently on these basic evaluations across all tasks.
Figure 26: Standard Prompting results for Sports-Athlete-Editing, across all localizations.
Figure 27: Standard Prompting results for Full-Sports-Editing, across all localizations.
Figure 28: Standard Prompting results for CounterFact-Editing, across all localizations.
Figure 29: Standard Prompting results for Sequential-CounterFact-Editing, across all localizations.
A.7.2 Details on Adversarial Prompt Evaluations
We report the adversarial prompt evaluations from Section 3.1 across all localizations. Figures 30, 31, 32 and 33 all show that FLU localization is more robust in MCQ compared to every other localization (significantly stronger for CounterFact). Figures 32 and 33 show that FLU localization is optimal in Paraphrase and Neighborhood in all cases except for Paraphrase compared to the Random localization in Sequential-CounterFact-Editing.
Figure 30: Adversarial Prompting results for Sports-Athlete-Editing, across all localizations.
Figure 31: Adversarial Prompting results for Full-Sports-Editing, across all localizations.
Figure 32: Adversarial Prompting results for CounterFact-Editing, across all localizations.
Figure 33: Adversarial Prompting results for Sequential-CounterFact-Editing, across all localizations.
A.7.3 Details on Adversarial Relearning
We retrain the model for 20 iterations with cross-entropy on half of the forget set (along with a standard retain and SFT loss), adding up all losses with loss coefficient 1.
We present relearning results for all localizations averaged over models. As shown in Figure 34, the FLU localization remains optimal, although the baselines of Nonlocalized, All MLPs, Random, and Random MLPs are competitive.
Figure 34: Relearning results for Sports-Athlete-Editing, across all localizations.
We also present relearning results on the other tasks. As mentioned in Section 3.2, since Full-Sports-Editing forget facts are not independent, we don’t expect valid results from relearning. Thus, in Figure 35, we see that every localization regains 100% editing accuracy.
Figure 35: Relearning results for Full-Sports-Editing.
On CounterFact-Editing and Sequential-CounterFact-Editing, as shown in Figures 36 and 37, none of the localizations relearn more than 7% accuracy, suggesting adversarial relearning was not a sufficiently strong enough evaluation for these tasks. Regardless, FLU localization is either the most or second-most robust localization to relearning, although localizations don’t differ by much.
Figure 36: Relearning results for CounterFact-Editing.
Figure 37: Relearning results for Sequential-CounterFact-Editing.
A.7.4 Details on Latent Knowledge
We present the probing classification accuracies for the three models separately here, as well as for all localizations we previously left out.
Figure 38: Linear probes applied to the forget set across all models, classifying model activations after various layers.
Figure 39: Linear probes applied to the forget set on Gemma-7B with 28 layers.
Figure 40: Linear probes applied to the forget set on Gemma-2-9b with 42 layers.
Figure 41: Linear probes applied to the forget set on Llama-3-8b with 32 layers.
In Gemma-7b and Llama-3-8b, FLU probing is the most monotonic and the best in the early layers, either steadily decreasing to 0 for forget accuracy or increasing to 1 for edit accuracy, with the least extreme peaks. In Gemma-2-9b, the Nonlocalized, All MLPs, and Random MLPs baselines are competitive with FLU. The other OT localization, Attribution Patching, has 100% probing forget accuracy across many layers, suggesting it represents the ground truth answer very clearly.
A.8 Soft Prompt Evaluations
Because many localizations seem to be weak to prompting schemes, we attempt a simple adaptive attack of soft prompts, where we optimize the continuous embeddings at the end of the prompt to recover the correct answer on half of our forget set. We then evaluate the model’s performance on the other half, with this soft prompt in place (Lester et al., 2021). We average over evaluations from four soft prompts. Soft prompt evaluations can be considered to be a more narrow form of few-shot finetuning, that is closer to searching for prompts that recover the model’s knowledge.
We find limited soft prompt success: across most tasks and models, we don’t recover much held-out forget set accuracy. On the Sports-Athlete-Editing, CounterFact-Editing, and Sequential-CounterFact-Editing tasks, Figures 42, 44 and 45 show that all localizations don’t significantly improve in Forget Accuracy over random chance, or are about equal between localizations, after soft prompts are applied. In Figure 43, specifically for Gemma-2 on Sports-Athlete-Editing we see some reasonable results with softprompts that are able to recover over 60% Forget Accuracy on OT localizations, while FLU, Nonlocalized, and All MLPs remain under 40% Forget Accuracy.
Figure 42: Metrics with soft prompts applied for Sports-Athlete-Editing, averaged over all models.
Figure 43: Metrics with soft prompts applied for Sports-Athlete-Editing for Gemma-2-9b.
Figure 44: Metrics with soft prompts applied for CounterFact-Editing, averaged over all models.
Figure 45: Metrics with soft prompts applied for Sequential-CounterFact-Editing, averaged over all models.